# Hyperspherical Latents Improve Continuous-Token Autoregressive Generation

### v1

@@ 4.2 Image Generation Result
VQGAN (esser2021taming, ) AR raster 1.4B 240 5.20 280.3 - -
ViT-VQGAN (yu2021vector, ) AR raster 1.7B 240 3.04 227.4 - -
LlamaGen-L (sun2024autoregressive, ) AR raster 343M 300 3.07 256.1 0.83 0.52
LlamaGen-XL* (sun2024autoregressive, ) AR raster 775M 300 2.62 244.1 0.80 0.57
LlamaGen-XXL* (sun2024autoregressive, ) AR raster 1.4B 300 2.34 253.9 0.80 0.59
RandAR-L (pang2025randar, ) AR random 343M 300 2.55 288.8 0.81 0.58
RandAR-XL (pang2025randar, ) AR random 775M 300 2.22 314.2 0.80 0.60
RandAR-XXL (pang2025randar, ) AR random 1.4B 300 2.15 322.0 0.79 0.62
RAR-B (yu2024randomized, ) AR hybrid 261M 400 1.95 290.5 0.82 0.58
RAR-L (yu2024randomized, ) AR hybrid 461M 400 1.70 299.5 0.81 0.60
MaskGIT (chang2022maskgit, ) Mask. random 227M 300 4.02 355.6 0.78 0.50
MAGVIT-v2 (yu2023language, ) Mask. random 307M 270 1.78 319.4 - -
VAR-d20 (tian2024visual, ) N.S. - 600M 350 2.57 302.6 0.83 0.56
VAR-d30 (tian2024visual, ) N.S. - 2B 350 1.92 323.1 0.82 0.59
@@
LDM-4 (rombach2022high, ) Diff. - 400M - 3.60 247.7 0.87 0.48
DiT-XL/2 (peebles2023scalable, ) Diff. - 675M 400 2.27 278.2 0.83 0.57
SiT-XL/2 (ma2024sit, ) Diff. - 675M 400 2.06 277.5 0.83 0.59
GIVT (tschannen2024givt, ) AR raster 1.67B 500 2.59 - 0.81 0.57
LatentLM-L (sun2024multimodal, ) AR raster 479M 400 2.24 253.8 - -
MAR-B (li2024autoregressive, ) Mask. random 208M 800 2.31 281.7 0.82 0.57
MAR-L (li2024autoregressive, ) Mask. random 479M 800 1.78 296.0 0.81 0.60
MAR-H (li2024autoregressive, ) Mask. random 943M 800 1.55 303.7 0.81 0.62
@@
We report Fréchet Inception Distance (FID) (heusel2017gans, ) as the primary metric, computed on 50k samples drawn with a fixed random seed using the ADM evaluation code (dhariwal2021diffusion, ). The optimal CFG scale is determined through a sweep with a step size of $0.1$. Following MAR (li2024autoregressive, ), we additionally report Inception Score (IS) (salimans2016improved, ) and Precision/Recall (Pre./Rec.) (kynkaanniemi2019improved, ).
@@
Overall, SphereAR delivers a scale-invariant AR model that sets the best reported FID with far fewer parameters and outperforms diffusion, masked-generation, and next-scale baselines. Appendix E shows qualitative results.
@@ 4.3 Ablation Study
The core design of SphereAR is to constrain latents on a hypersphere via S-VAE. We compare it with prior VAEs. In particular, we evaluate three settings: (1) Diagonal-Gaussian ($\beta$-VAE). A standard diagonal-Gaussian posterior trained with an up-weighted KL term. [^2] Most prior VAEs compute the KL by taking a sum over spatial-channel dimensions $(h\times w\times d)$ and then a batch mean. We instead take a mean over all elements (batch and spatial-channel), which lowers the numerical KL value; to match the effective regularization strength, we therefore use larger KL weights (our $10^{-2}$ roughly matches a prior $2\times 10^{-6}$). We sweep four KL weights $\{0.01,\,0.04,\,0.08,\,0.16\}$, denoted G-01, G-04, G-08, and G-16. (2) $\sigma$-VAE (fixed variance). Following LatentLM (sun2024multimodal, ), we fix the posterior scale with a fixed, non-learned scalar $\sigma\sim\mathcal{N}(0,\,C_{\sigma})$, and sample $\mathbf{z}=\bm{\mu}_{\phi}(\mathbf{x})+\sigma\,\bm{\epsilon}$ with $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We sweep $C_{\sigma}\in\{0.1,\,0.2,\,0.5\}$ (denoted F-01, F-02, F-05). (3) S-VAE (hyperspherical). A Power Spherical posterior on $\mathbb{S}^{d-1}$ with a KL-weight sweep $\{0.001,\,0.004,\,0.008\}$, denoted S-01, S-04, S-08.
@@
Gaussian 2.97 240.2
Gaussian 2.89 254.3
Gaussian 2.68 257.3
Spherical 2.52 258.4
The above ablations indicate that both the normalization on latent tokens and the hyperspherical posterior are important. Because the normalization can affect two interfaces—the VAE decoder’s input and the AR model’s inputs/outputs—we further isolate their effects by conducting a variant that normalizes only the VAE decoder’s input. As summarized in Table 2, normalization applied to the AR inputs and outputs is more critical: normalizing only the VAE decoder’s input yields a modest gain (FID $2.97\!\to\!2.89$), whereas additionally normalizing AR inputs/outputs produces a larger improvement (FID $2.89\!\to\!2.68$). This matches our analysis: the AR pathway re-feeds tokens step by step, so scale errors would otherwise accumulate, while the VAE decoder consumes its input once and does not induce cascading scale drift. Finally, replacing the diagonal-Gaussian posterior with a hyperspherical one gives a further boost (FID $2.68\!\to\!2.52$), confirming that aligning the posterior with constant-norm geometry is beneficial.
@@ 5 Conclusion
While our results substantiate the motivation and design choices of SphereAR, several extensions would further strengthen this work: (i) exploring Riemannian Flow Matching (RFM) (chen2023flow, ), which may better align with hyperspherical latent geometry since trajectories of RFM remain on the hypersphere; and (ii) extending SphereAR to multimodal applications. We leave these to future work.
Acknowledgments
We are grateful to Prof. Di He for his careful reading and helpful comments on earlier drafts of this manuscript.
@@ Appendix C Comparison with MAR’s VAE
We perform an ablation that swaps our S-VAE for MAR’s VAE [12] while keeping the AR backbone, training schedule, and evaluation protocol identical (see Sec. 4.3). As summarized in Table 3, S-VAE yields a large FID gain (4.54 $\rightarrow$ 2.52) and higher IS (241.6 $\rightarrow$ 258.4). These results indicate that constant-norm, directional latents from S-VAE materially strengthen continuous-token AR generation.
@@
Most latent VAEs for image generation adopt a VQGAN-style encoder-decoder [3], i.e., a convolutional (CNN) backbone with downsampling/upsampling blocks. This design is parameter-efficient ($\sim$70M) but throughput is often limited by activation memory and bandwidth on large feature maps, leading to slow training and inference. ViT-VQGAN [4] replaces the CNN backbone with a Vision Transformer (ViT) to improve efficiency; however, as shown in Table 4, a pure ViT backbone yields weaker generative metrics than a CNN.
@@
In our implementation, to match the parameter scale of a VQGAN-style CNN encoder-decoder [3], our S-VAE uses 6 ViT blocks in the encoder and 12 in the decoder, each with hidden size 512. The encoder’s CNN stem performs patchification via 4 downsampling stages (overall $16\times$ reduction) with channel widths $[64,64,128,256,512]$; the decoder mirrors this with 4 upsampling stages and an extra residual block per stage. In total, the S-VAE has $\sim$75M parameters.
We compare three encoder-decoder backbones under the same training setup and losses: (i) a VQGAN-style CNN [3]; (ii) a pure ViT [4] (12 Transformer blocks in both encoder and decoder, hidden size 768; $\sim$170M params); and (iii) our Hybrid design. To reduce compute, the training follows the setting in Sec. 4.3. We report training throughput (iterations/s) of VAE, perceptual distortion (LPIPS with VGG-16), and downstream ImageNet generative metrics. Results in Table 4 show: ViT is fastest (7.19 it/s) but slightly worse on LPIPS/FID/IS; the CNN is slowest (2.25 it/s) yet competitive in IS; our Hybrid retains the best reconstruction (LPIPS 0.166) and the best FID (2.52) while running at 5.81 it/s—about $2.6\times$ faster than the CNN and at $81\%$ of ViT throughput (5.81 vs 7.19 it/s). Overall, the hybrid backbone offers the most favorable speed-quality trade-off.
@@
Appendix E Model Generated Examples

### v2

@@ 4.2 Image Generation Result
VQGAN (Esser et al., 2021) AR raster 1.4B 240 5.20 280.3 - -
ViT-VQGAN (Yu et al., 2021) AR raster 1.7B 240 3.04 227.4 - -
LlamaGen-L (Sun et al., 2024a) AR raster 343M 300 3.07 256.1 0.83 0.52
LlamaGen-XL* (Sun et al., 2024a) AR raster 775M 300 2.62 244.1 0.80 0.57
RandAR-L (Pang et al., 2025) AR random 343M 300 2.55 288.8 0.81 0.58
RandAR-XL (Pang et al., 2025) AR random 775M 300 2.22 314.2 0.80 0.60
RAR-B (Yu et al., 2024) AR hybrid 261M 400 1.95 290.5 0.82 0.58
RAR-L (Yu et al., 2024) AR hybrid 461M 400 1.70 299.5 0.81 0.60
MaskGIT (Chang et al., 2022) Mask. random 227M 300 4.02 355.6 0.78 0.50
MAGVIT-v2 (Yu et al., 2023) Mask. random 307M 270 1.78 319.4 - -
VAR-d20 (Tian et al., 2024) N.S. - 600M 350 2.57 302.6 0.83 0.56
VAR-d30 (Tian et al., 2024) N.S. - 2B 350 1.92 323.1 0.82 0.59
@@
LDM-4 (Rombach et al., 2022) Diff. - 400M - 3.60 247.7 0.87 0.48
DiT-XL/2 (Peebles and Xie, 2023) Diff. - 675M 400 2.27 278.2 0.83 0.57
SiT-XL/2 (Ma et al., 2024) Diff. - 675M 400 2.06 277.5 0.83 0.59
GIVT (Tschannen et al., 2024) AR raster 1.67B 500 2.59 - 0.81 0.57
LatentLM-L (Sun et al., 2024b) AR raster 479M 400 2.24 253.8 - -
MAR-B (Li et al., 2024a) Mask. random 208M 800 2.31 281.7 0.82 0.57
MAR-L (Li et al., 2024a) Mask. random 479M 800 1.78 296.0 0.81 0.60
MAR-H (Li et al., 2024a) Mask. random 943M 800 1.55 303.7 0.81 0.62
@@
We report Fréchet Inception Distance (FID) (Heusel et al., 2017) as the primary metric, computed on 50k samples drawn with a fixed random seed using the ADM evaluation code (Dhariwal and Nichol, 2021). The optimal CFG scale is determined through a sweep with a step size of $0.1$. Following MAR (Li et al., 2024a), we additionally report Inception Score (IS) (Salimans et al., 2016) and Precision/Recall (Pre./Rec.) (Kynkäänniemi et al., 2019).
@@
Overall, SphereAR delivers a scale-invariant AR model that sets the best reported FID with far fewer parameters and outperforms diffusion, masked-generation, and next-scale baselines. Appendix K shows qualitative results.
@@ 4.3 Ablation Study
The core design of SphereAR is to constrain latents on a hypersphere via S-VAE. We compare it with prior VAEs. In particular, we evaluate three settings: (1) Diagonal-Gaussian ($\beta$-VAE). A standard diagonal-Gaussian posterior trained with an up-weighted KL term. [^2] Most prior VAEs compute the KL by taking a sum over spatial-channel dimensions $(h\times w\times d)$ and then a batch mean. We instead take a mean over all elements (batch and spatial-channel), which lowers the numerical KL value; to match the effective regularization strength, we therefore use larger KL weights (our $10^{-2}$ roughly matches a prior $2\times 10^{-6}$). We sweep four KL weights $\{0.01,\,0.04,\,0.08,\,0.16\}$, denoted G-01, G-04, G-08, and G-16. (2) $\sigma$-VAE (fixed variance). Following LatentLM (Sun et al., 2024b), we fix the posterior scale with a fixed, non-learned scalar $\sigma\sim\mathcal{N}(0,\,C_{\sigma})$, and sample $\mathbf{z}=\bm{\mu}_{\phi}(\mathbf{x})+\sigma\,\bm{\epsilon}$ with $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We sweep $C_{\sigma}\in\{0.1,\,0.2,\,0.5\}$ (denoted F-01, F-02, F-05). (3) S-VAE (hyperspherical). A Power Spherical posterior on $\mathbb{S}^{d-1}$ with a KL-weight sweep $\{0.001,\,0.004,\,0.008\}$, denoted S-01, S-04, S-08.
@@
1 Gaussian 2.97 240.2
2 Gaussian 2.89 254.3
3 Gaussian 2.68 257.3
4 Spherical 2.52 258.4
The above ablations indicate that both the normalization on latent tokens and the hyperspherical posterior are important. Because the normalization can affect two interfaces—the VAE decoder’s input and the AR model’s inputs/outputs—we further isolate their effects by conducting a variant that normalizes only the VAE decoder’s input. As summarized in Table 2, normalization applied to the AR inputs and outputs is more critical. In particular, normalizing only the VAE decoder’s input yields a modest gain (FID $2.97\!\to\!2.89$, compare No. 1 with No. 2); whereas additionally normalizing AR inputs/outputs produces a larger improvement (FID $2.89\!\to\!2.68$, compare No.2 with No. 3). This finding supports our analysis: the AR pathway re-feeds tokens step by step, so scale errors would otherwise accumulate. In contrast, the VAE decoder consumes its input in a single pass and does not induce such cascading scale drift. Finally, adopting the Spherical posterior (No. 4) provides a further significant boost (FID $2.68\!\to\!2.52$), confirming that aligning the posterior with constant-norm geometry is beneficial.
@@ 5 Conclusion
While our results substantiate the motivation and design choices of SphereAR, several extensions would further strengthen this work: (i) exploring Riemannian Flow Matching (RFM) (Chen and Lipman, 2023), which may better align with hyperspherical latent geometry since trajectories of RFM remain on the hypersphere; and (ii) extending SphereAR to more datasets and multimodal applications. We leave these to future work.
@@ Appendix C Comparison with MAR’s VAE
We perform an ablation that swaps our S-VAE for MAR’s VAE [Li et al., 2024a] while keeping the AR backbone, training schedule, and evaluation protocol identical (see Sec. 4.3). As summarized in Table 3, S-VAE yields a large FID gain (4.54 $\rightarrow$ 2.52) and higher IS (241.6 $\rightarrow$ 258.4). These results indicate that constant-norm, directional latents from S-VAE materially strengthen continuous-token AR generation.
@@
Most latent VAEs for image generation adopt a VQGAN-style encoder-decoder [Esser et al., 2021], i.e., a convolutional (CNN) backbone with downsampling/upsampling blocks. This design is parameter-efficient ($\sim$70M) but throughput is often limited by activation memory and bandwidth on large feature maps, leading to slow training and inference. ViT-VQGAN [Yu et al., 2021] replaces the CNN backbone with a Vision Transformer (ViT) to improve efficiency; however, as shown in Table 4, a pure ViT backbone yields weaker generative metrics than a CNN.
@@
In our implementation, to match the parameter scale of a VQGAN-style CNN encoder-decoder [Esser et al., 2021], our S-VAE uses 6 ViT blocks in the encoder and 12 in the decoder, each with hidden size 512. The encoder’s CNN stem performs patchification via 4 downsampling stages (overall $16\times$ reduction) with channel widths $[64,64,128,256,512]$; the decoder mirrors this with 4 upsampling stages and an extra residual block per stage. In total, the S-VAE has $\sim$75M parameters.
We compare three encoder-decoder backbones under the same training setup and losses: (i) a VQGAN-style CNN [Esser et al., 2021]; (ii) a pure ViT [Yu et al., 2021] (12 Transformer blocks in both encoder and decoder, hidden size 768; $\sim$170M params); and (iii) our Hybrid design. To reduce compute, the training follows the setting in Sec. 4.3. We report training throughput (iterations/s) of VAE, perceptual distortion (LPIPS with VGG-16), and downstream ImageNet generative metrics. Results in Table 4 show: ViT is fastest (7.19 it/s) but slightly worse on LPIPS/FID/IS; the CNN is slowest (2.25 it/s) yet competitive in IS; our Hybrid retains the best reconstruction (LPIPS 0.166) and the best FID (2.52) while running at 5.81 it/s—about $2.6\times$ faster than the CNN and at $81\%$ of ViT throughput (5.81 vs 7.19 it/s). Overall, the hybrid backbone offers the most favorable speed-quality trade-off.
@@
Appendix E ImageNet $\times512$
Table 5: Overall comparison on ImageNet $\times512$ class-conditional generation.
[…]
Discrete Tokens
MaskGIT [Chang et al., 2022] Mask. random 227M 300 4.46 342.0
MAGVIT-v2 [Yu et al., 2023] Mask. random 307M 270 1.91 324.3
VAR-d36-s [Tian et al., 2024] N.S. - 2.3B 350 2.62 303.2
Continuous Tokens
DiT-XL/2 [Peebles and Xie, 2023] Diff. - 675M 400 3.04 240.8
MAR-L [Li et al., 2024a] Mask. random 481M 800 1.73 279.9
SphereAR-L (Our) AR raster 481M 400 1.67 292.5
Following prior work, we also evaluate on ImageNet at $512\times 512$ resolution. We train S-VAE at $512\times 512$ with hidden dimension $d=16$ and a KL coefficient $\beta=0.016$. For the autoregressive transformer, we adopt the SphereAR-L configuration and set the number of class-conditional tokens to 64. Unless otherwise noted, all other settings mirror the $256\times 256$ experiments. As shown in Table 5, SphereAR-L achieves an FID of 1.67 competitive with prior models.
Appendix F Training Convergence Efficiency
Figure 6: Training Convergence Efficiency Comparison. We plot the generative performance (FID, lower is better) as a function of training cost (wall-clock time) for SphereAR (our model), MAR, and VAR. SphereAR converges to a strong FID score significantly faster than both baselines.
One key advantage of autoregressive (AR) models is their high training efficiency. More specifically, for a single training sample, an AR model calculates a loss across all positions of the sequence (i.e., all prediction steps) in one forward pass (using teacher forcing). In contrast, diffusion or masked-based models merely compute a loss for a single sampled step (e.g., one timestep t in diffusion or one mask pattern). Therefore, because diffusion or masked-based models receive a less dense supervisory signal per sample, they often require significantly more epochs to converge.
To explicitly demonstrate this, we additionally evaluated the generative performance of intermediate checkpoints during the training of SphereAR. For a fair comparison, all wall-clock training times were measured using 8 NVIDIA A100 GPUs. As shown in Fig. 6, SphereAR converges much faster than MAR and VAR. This efficiency is stark: in terms of epochs, SphereAR (200 epochs) achieves comparable performance to MAR-L (800 epochs). In terms of wall-clock time, SphereAR achieves this same performance using only $\sim$20% of the training cost required by MAR-L.
Appendix G Inference Speed
Figure 7: Inference speed comparison on a single A100 GPU. SphereAR’s speed, which depends on the number of diffusion steps, is benchmarked against MAR (slowest), VAR (fastest), and LlamaGen.
This section benchmarks the inference speed of our model against previous approaches. All benchmarks were performed on a single NVIDIA A100 GPU across various batch sizes. The results, presented in Fig. 7, lead to the following observations: (i) VAR is the fastest model, as it only requires dozens of steps to generate an image. The purely autoregressive model LlamaGen is also highly efficient. (ii) MAR is the slowest. This is expected, as its architecture requires accessing the full-length token sequence at each sampling step, which is compounded by an additional 100 sampling steps in its diffusion head. (iii) For SphereAR, our default configuration (with 100 sampling steps in diffusion head) is significantly faster than MAR but slower than LlamaGen due to the computational overhead of the diffusion head. (iv) We also evaluated two additional SphereAR settings (with 10 and 1 diffusion steps, respectively), which exhibit progressively faster speeds. This result clearly demonstrates that the primary inference bottleneck lies in the diffusion head.
These observations highlight several paths for optimization. First, the high sampling cost of the diffusion head can be mitigated by integrating recent few-step diffusion models, such as Shortcut-Model [Frans et al., 2024] or MeanFlow [Geng et al., 2025], which dramatically reduce the required number of steps. Second, since the core of our model is a standard autoregressive Transformer (akin to other LLMs), it can directly benefit from established inference optimization libraries (e.g., vLLM) to further accelerate the token generation process.
Appendix H Robustness with high CFG scales
Figure 8: Performance (FID, lower is better) versus CFG scale, comparing our S-VAE model (‘S-04‘) with the Gaussian baseline (‘G-16‘). The curves demonstrate that ‘S-04‘ is more robust to high CFG values.
To demonstrate SphereAR’s robustness to high CFG scales, we conducted an additional experiment. In particular, we ran sampling with progressively higher CFG scales for the best diagonal-Gaussian model (G-16) and our best S-VAE model (S-04) from Sec. 4.3. The results are presented in Fig. 8.
As shown in the figure, the performance of both models’ degrades as the CFG scale increases. However, the rate of degradation for S-04 is much slower. This is evidenced by the widening performance gap: at a CFG scale of 4.0, the gap between the models is $\sim$1.0, but at a CFG scale of 8.0, it widens to $\sim$2.0. This result clearly demonstrates that the proposed model is more robust to aggressive CFG scales.
Appendix I Failure Cases and Limitations
We present a qualitative analysis of failure modes in Fig. 9(b). Similar to existing methods, SphereAR can produce results with noticeable artifacts, especially in complex classes. However, even in these failure cases, SphereAR-L consistently produces visually superior results with fewer artifacts than the baseline MAR-L.
Despite this relative improvement, both models share common limitations. These primarily include artifacts in human faces and, most prominently, a failure to render legible text or other high-frequency details (e.g., street signs, bottle labels), which are often generated as semantically meaningless patterns.
We posit this is largely a data-driven limitation. As noted by MAR [Li et al., 2024a], models trained on academic datasets like ImageNet often show a "noticeable gap in visual quality" compared to commercial models trained on massive, diverse web data, especially for categories like coherent text and faces. These failures are thus likely attributable to the dataset’s limitations rather than a fundamental flaw in our architecture.
Due to computational constraints, our current evaluation is focused on the ImageNet benchmark. Therefore, assessing the scalability of SphereAR and its robustness in more diverse, large-scale data regimes remains a critical direction for future investigation.
(a) MAR-L
(b) SphereAR-L
Figure 9: Failure cases. Similar to existing methods, our model can produce results with noticeable artifacts. We show MAR-L (left) and SphereAR-L (right) ’s results of the same classes.
Appendix J Use of LLMs
We used a large language model (OpenAI’s ChatGPT) solely for language polishing—proofreading, minor rephrasing, and tone consistency. It did not generate or influence the scientific ideas, methods, results, or conclusions. All suggestions were reviewed and edited by the authors, who take full responsibility for the paper’s content and integrity.
Appendix K Model Generated Examples